OpenStudy (anonymous):

5% of (35 x 102) = _____?

OpenStudy (inkyvoyd):

178.5

OpenStudy (apoorvk):

that is. 0.05 x (3570) = 178.5

OpenStudy (anonymous):

\[5\%\text{ of }(35\times102)\] \[=\frac{5}{100}\times(35 \times 102)\]

OpenStudy (anonymous):

thank you